# AI-Powered Episodic Future Thinking

### v1

@@ 6 Limitations and Future work
Our study design only measured immediate user perceptions and experiences. We did not track longer-term behavioral or psychological outcomes. In subsequent research, the research team aims to conduct longer-term studies to observe how continued use of EFTeacher’s generated cues might influence users’ behavior over time. This longitudinal approach will allow us to assess whether the chatbot’s interactive EFT features have a sustained impact on users’ motivation, goal-setting, and health-related behaviors.
In parallel, future research should compare user perceptions and the efficacy of EFTeacher to the existing survey method and clinician-led interviews. Although the survey method was designed as a more scalable alternative to the human interview, poor cue quality cues in this self-administered method (e.g., vague or non-episodic) may reduce treatment fidelity and thus efficacy of the intervention. Although EFT is a promising behavioral intervention to facilitate weight loss and dietary [17, 18, 19, 20, 21], significant effects on health behaviors are not always observed [20, 19]. Thus, future research should examine whether the method used to generate cues moderates these behavioral effects and user perceptions.
Despite its limitations, our study offers initial evidence supporting the feasibility and promise of a chatbot-based EFT intervention, suitable to guide follow-on studies.
7 Conclusion
@@
In this study, we evaluate EFTeacher, an AI chatbot built on GPT-4-Turbo, designed to assist participants in identifying meaningful future events and creating detailed text descriptions (cues) for those events. We investigated users’ overall experiences with EFTeacher, their impressions of the chatbot, and their perceptions of working with an interactive AI tool for EFT generation.
@@
Our findings highlighted both strengths and areas for improvement in EFTeacher’s design and functionality. Quantitative analysis of user impressions, represented through content charectristics of vividness, liking, excitement, and importance ratings, revealed positive trends.
@@
Analysis of user feedback revealed five key themes that provided deeper insights into users’ experiences with EFTeacher. Participants appreciated the chatbot’s vivid and descriptive language, which enhanced engagement and encouraged meaningful self-reflection. This feature helped users articulate their goals and vividly visualize future scenarios. However, some participants found the detailed language excessive, emphasizing the need for adaptability in response style. The chatbot’s adaptability features—enabling refinements in tone and style—were identified as a significant strength. However, these features were underutilized by some participants, indicating the need for better user guidance. Participants also highlighted challenges with repetitive task structures. Reducing repetitive patterns could improve usability and promote a more natural conversational experience.
@@
Overall, EFTeacher demonstrated potential as an effective tool for guiding users through EFT generation. Its responses were highly valued, with participants emphasizing its ability to inspire reflection on their goals. The study identified areas for improvement, such as minimizing repetitive patterns and offering clearer guidance on adaptability features. Our design suggestions could further enhance the application of AI chatbots based on LLMs in EFT interventions for maladaptive health behaviors.
Acknowledgment

### v2

@@ 5 Limitations, Conclusions, and Follow-on Work
Limitations: Our study design only measured immediate user perceptions and experiences. We did not track longer-term behavioral or psychological outcomes. In subsequent research, the research team aims to conduct longer-term studies to observe how continued use of EFTeacher’s generated cues might influence users’ behavior over time. This longitudinal approach would allow us to assess whether the chatbot’s interactive EFT features have a sustained impact on users’ motivation, goal-setting, and health-related behaviors. Despite its limitations, our study offers initial evidence supporting the feasibility and promise of a chatbot-based EFT intervention, suitable to guide follow-on studies. Our design suggestions could further enhance the application of AI chatbots based on LLMs in EFT interventions for maladaptive health behaviors.
@@
Conclusions: In this study, we evaluated EFTeacher, an AI chatbot built on GPT-4-Turbo, designed to assist participants in identifying meaningful future events and in creating detailed text descriptions (cues) for those events. We investigated users’ usability and feasibility experiences with EFTeacher, their impressions of the chatbot, and their perceptions of working with an interactive AI tool for EFT generation.
@@
Our findings highlighted both strengths and areas for improvement in EFTeacher’s design and functionality. Analysis of user impressions, represented through content characteristics of vividness, liking, excitement, and importance ratings, revealed positive trends.
@@
Analysis of user feedback revealed five key themes that provided deeper insights into users’ experiences with EFTeacher. Participants appreciated the chatbot’s vivid and descriptive language, which enhanced engagement and encouraged self-reflection. This feature helped users articulate their goals and vividly visualize future scenarios. However, some participants found the detailed language excessive, emphasizing the need for adaptability in response style. The chatbot’s adaptability features—enabling refinements in tone and style—were identified as a significant strength. However, these features were underutilized by some participants, indicating the need for better user guidance. Participants also highlighted challenges with repetitive task structures. Reducing repetitive patterns could improve usability and promote a more natural conversational experience.
@@
Overall, EFTeacher demonstrated potential as a tool for guiding users through EFT generation. Its responses were valued, with participants emphasizing its ability to facilitate imagination and inspire reflection on their goals. The study identified areas for improvement, such as minimizing repetitive patterns and offering clearer guidance on adaptability features.
Based on feedback from our qualitative research discussed, we refined the chatbot to generate EFT cues for four timeframes (events) to reduce repetition and enhance user experience and engagement, allowing users to discuss events occurring in one month, three months, one year, and three years.
Follow-on Work: As a follow-on to the study reported above, we designed a second study, with more participants. The aims were to: further test our hypotheses, scale up our investigation, and compare EFTeacher with existing alternatives, i.e., to the survey method, or to clinician-led interviews. Since it was essential to test if EFTeacher led to cues comparable to those from the other methods, and that those cues would reduce delay discounting at least as well as the other methods, we added a final phase to the user session to assess effects on delay discounting. Since that lengthened the time required of participants, and since we already had feedback and had made changes suggested from open-ended questions, we removed the open-ended questions and reduced the number of generated cues from seven to four.
Preliminary analysis of the second study indicates that EFTeacher helps participants generate cues that are comparable to those of the two other methods. They also suggest that there is as good an effect on delay discounting as with the other methods, so that at least as good an effect on health outcomes would result from use of EFTeacher. We plan on subsequent reports of our findings, once we have completed the analysis and comparison of the methods in our second study, the cue texts generated, their reception by participants, and the effect on delay discounting.
Acknowledgments